# A Survey of 3D Reconstruction with Event Cameras: From Event-based Geometry to Neural 3D Rendering

### v1

@@ 2.3 Event-based 3D Reconstruction Types
Schraml et al.[85] 2015 Trajectory Event accumulation frame 3D depth map (✗) -
Zhu et al.[119] 2018 Velocity Event disparity volume 3D depth map (✓) MVSEC
Zhou et al.[117] 2018 Pose Time surface 3D depth map (✗) MVSEC
@@
Leng et al.[38] 2018 Pose Time surface Point cloud (✗) -
Steffen et al.[87] 2019 - 4D vector Voxel (✗) -
@@
Zhou et al.[118] 2021 Pose Time surface Point cloud, depth map (✗) RPG, ESIM
Nam et al.[62] 2022 Pose Event Stack 3D depth map (✓) DSEC
@@
Ghosh et al.[29] 2022 Pose Event disparity volume 3D depth map (✗) MVSEC, DSEC, ESIM
Table 3: Monocular event camera: Geometry-based methods with traditional outputs
Author Year Priors Event Rep. Output (Dense?) Dataset
Kim et al.[41] 2016 Trajectory Event gradient accumu. 3D depth map (✗) -
Rebecq et al.[77] 2016 Pose Event frame 3D depth map (✗) Multi-Keyframe
Rebecq et al.[75] 2018 Trajectory Event-by-event 3D depth map (✗) -
Guan et al.[30] 2024 Trajectory Time surface 3D depth map (✓) DAVIS240c
Elms et al.[17] 2024 Trajectory Time surface Sparse point cloud (✗) TOPSPIN
Table 4: Monocular event camera: Deep learning-based methods with traditional outputs
Author Year Priors Pipeline Event Rep. Model Output Dataset
Baudron et al.[3] 2020 Silhouette Event accumu. frames E2S(CNN) Mesh ShapeNet
Xiao et al.[100] 2022 Pose Voxel grid E2VID(RNN-CNN) Mesh ESIM
Wang et al.[100] 2023 Contour, Trajec. Voxel grid Evac3d(CNN) Mesh MOEC-3D
Chen et al.[12] 2023 - Event frame E2V(CNN) Voxel SynthEVox3D
Xu et al. [102] 2025 - Event frame (sobel) E2V(CNN) Voxel SynthEVox3D
Table 5: Monocular event camera: NeRF-based Methods
Author Model Yr-Mo Inputs Event Rep. Colourful Dataset
Klenk et al. [43] E-NeRF 2023-01 Event stream EAF ESIM
Hwang et al. [37] Ev-NeRF 2023-03 Event stream EAF IJRR, HQF
Rudnev et al. [81] EventNeRF 2023-06 RGB bayer event stream EAF NeRF dataset
Low et al. [54] Robust e-NeRF 2023-10 Event stream Event-by-event TUM-VIE
Bhattacharya et al. [4] EvDNeRF 2024-01 Event stream EAF Real-Fork
Wang et al. [92] NeRF(Enhanced) 2024-05 Event stream EAF PAEv3D
Feng et al. [19] AE-NeRF 2025-01 Event stream Event-by-event TUM-VIE
Table 6: Monocular event camera: 3DGS-based Methods
Author Model Yr-Mo Inputs Event Rep. Colourful Dataset
Wang et al. [93] EvGGS 2024-05 Event stream Voxel Grid, EAF Ev3DS
Zhang et al. [114] Elite-EvGS 2024-09 Event stream EAF EventNeRF dataset
Wu et al. [97] Ev-GS 2024-09 Event stream EGA NeRF dataset
Huang et al. [35] IncEventGS 2024-10 Event stream EAF Replica dataset
Yin et al. [106] E-3DGS 2024-10 Event stream Event-by-event EventNeRF dataset
Yura et al. [110] EventSplat 2024-12 RGB bayer event stream EAF NeRF dataset
Zahid et al. [111] E-3DGS 2025-02 RGB bayer event stream EAF E-3DGS dataset
Table 7: Event cameras in multimodal system: Methods with traditional outputs
Author Year Device Priors Event Rep. Real-time Output (Dense?)
Leroux et al.[49] 2018 Structured light Pose Time surface Point cloud (✓)
Huang et al.[36] 2021 Structured light Pose Event-by-event Point cloud (✓)
Zuo et al.[121] 2022 D-RGB camera Trajectory Time surface Point cloud (✗)
Xiao et al.[99] 2023 Structured light Pose Event frame Point cloud (✓)
Fu et al.[20] 2023 Structured light Pose Time surface Point cloud (✓)
@@
Li et al.[50] 2024 Structured light Pose Event-by-event Point cloud (✓)
Table 8: Event cameras in multimodal system: NeRF-based methods
Author Model Yr-Mo Inputs Event Rep. Colourful Dataset
Qi et al. [69] E2NeRF 2023-10 Blurry RGB, Event stream EAF NeRF dataset
Ma et al. [56] Deformable Event-NeRF 2023-10 Blurry RGB, Event Stream Event-by-event HS-ERGB, CED, EVIMOv2
Cannici et al. [8] Ev-DeblurNeRF 2024-06 Blurry RGB, Event stream EAF Ev-DeblurBlender
Qi et al. [70] E3NeRF 2024-08 Blurry RGB, Event stream EGA Real-World-Challenge Dataset
Tang et al. [89] LSE-NeRF 2024-09 Blurry RGB, Event stream EAF LSE-NeRF, EVIMOv2
Qi et al. [71] EBAD-NeRF 2024-10 Blurry RGB, Event stream EAF Deblur-NeRF dataset
Chen et al. [14] Event-ID 2024-10 Blurry RGB, Event stream EAF NeRF synthetic dataset
Rudnev et al. [82] Dynamic EventNeRF 2024-12 Blurry RGB, Event stream EAF -
Table 9: Event cameras in multimodal system: 3DGS-based methods
Author Model Yr-Mo Inputs Event Rep. Colourful Dataset
Yu et al. [108] EvaGaussians 2024-05 Blurry RGB, Event stream EGA EvaGaussians
Xiong et al. [101] Event3DGS 2024-06 Blurry RGB, Event stream EGA EventNeRF
Weng et al. [96] EaDeblur-GS 2024-07 Blurry RGB, Event stream EGA E2NeRF
Liao et al. [51] EF-3DGS 2024-10 Blurry RGB, Event stream EAF Tanks and Temples
Deguchi et al. [15] E2GS 2024-10 Blurry RGB, Event stream EAF NeRF dataset
Xu et al. [103] EventBoosted-3DGS 2024-11 Blurry RGB, Event stream EGA -
Wu et al. [98] SweepEvGS 2024-12 Static RGB, Event stream EGA NeRF synthetic dataset
3 Methods with Stereo Event Cameras
Stereo event cameras typically refer to two or more rigidly mounted event cameras. A stereo vision system composed of stereo event cameras is capable of asynchronously capturing event streams from the left and right viewpoints of a scene, enabling event-driven stereo matching and 3D scene reconstruction. In general, the goal of stereo matching is to identify the most probable pairs of left and right events in both space and time, in order to compute disparity and subsequently estimate the depth structure of the scene.
Event-driven 3D-related tasks were initially pioneered using stereo event cameras in the 2010s [44, 68]. However, these early works do not achieve a fully modelled 3D reconstruction as the final output, and therefore are not included in our review.
Based on the data processing pipeline, 3D reconstruction methods using only stereo event cameras can be categorised into: (3.1) geometry-based methods and (3.2) learning-based methods. Table 2 and Figure 6 provide an overview of these methods.
Figure 6: Overview of methods with stereo event camera. (a) Schematic of general disparity matching (b) Refocused DSI Fusion to replace disparity matching module, from Ghosh et al. [29] (c) Self-Organising Maps from Steffen et al. [87] (d) UNet-ResNet (CNN) from Nam et al. [62]
3.1 Geometry-based Methods
With disparity matching: Many earlier approaches perform depth mapping reconstruction by computing disparity between events observed at the same timestamp across different viewpoints, followed by geometry-based multi-view stereo estimation to achieve real-time 3D depth reconstruction. In 2015, Schraml et al. [85] introduced a novel stereo matching method for a rotating panoramic stereo event system. Their approach defines a cost metric based on event distribution similarity between left and right views and computes disparity through dynamic programming on sparse event maps. The resulting disparities are then used to reconstruct real-world depth in 360° panoramic views. In 2018, Zhu et al.[119] proposed a method that synchronises events in time using known camera velocities, constructs a dense event disparity volume, and performs real-time sliding window matching, introducing a novel matching cost function combining ambiguity and similarity. Simultaneously, Leng et al.[38] reformulated event matching as a time-based stereo vision problem, where disparity estimation is achieved by detecting temporal coincidence of events across epipolar lines. Their method introduces a generalised spiking neuron model that integrates spatio-temporal event information with temporal decay, allowing robust matching under varying motion and illumination conditions. A voting scheme is then used to infer depth from the accumulated neural activations, enabling dense 3D reconstruction from pure event streams.
@@ 3.1 Geometry-based Methods
Without disparity matching: Disparity matching can also be bypassed altogether by directly optimising geometric or temporal consistency across event streams. In 2018, Zhou et al. [117] proposed a novel forward-projection based depth estimation method, which avoids the need for explicit disparity computation. Instead of finding correspondences between stereo events, their method optimises a temporal consistency energy by projecting candidate depth hypotheses into both stereo time surfaces and measuring their alignment. A coarse-to-fine search strategy is used to obtain semi-dense depth maps in real-time, enabling robust depth reconstruction under challenging conditions. Building upon this idea, Zhou et al. [118] extended the concept to a stereo event-based visual odometry framework. By jointly optimising over spatial and temporal consistency across stereo time surfaces, they estimate not only per-frame depth but also the stereo camera motion. The system runs in real-time and outputs semi-dense depth maps aligned with visual odometry estimates, making it well-suited for 3D reconstruction in dynamic environments. Later in 2022, Ghosh et al. [29] proposed a multi-event-camera fusion framework that leverages Disparity Space Images (DSIs) [45] to accumulate event ray densities from multiple viewpoints. Instead of requiring event correspondences, their method refocuses events in DSI space and performs probabilistic fusion to obtain high-quality depth maps. They further introduce an outlier rejection strategy to enhance robustness and demonstrate generalisation across multiple datasets without fine-tuning.
3.2 Learning-based Methods
@@
Deep learning processing: More recent methods leverage deep neural networks to learn stereo disparity estimation and event representations. In 2022, Nam et al. [62] introduced a deep stereo framework that combines multi-density event stacking with attention-guided encoding via a UNet-ResNet [80, 32] backbone. Their method further incorporates future event prediction during training through a distillation loss [34, 86], enabling real-time, dense depth estimation with high accuracy on the DSEC dataset [27].
4 Methods with Monocular Event Cameras
@@
Despite these limitations, monocular setups are more lightweight, power-efficient, and easier to deploy and move, making them highly suitable for embedded systems and mobile equipment. Consequently, monocular event-based 3D reconstruction has attracted increasing research interest and covers a broad range of tasks and 3D representation strategies.
Existing monocular approaches can be broadly divided into four categories based on their underlying modelling techniques and output formats: (4.1) Geometry-based methods with traditional outputs, (4.2) Learning-based methods with traditional outputs, (4.3) methods with Neural Radiance Fields, and (4.4) methods with 3D Gaussian Splatting. The first two categories aim to reconstruct explicit 3D outputs such as depth maps, point clouds, meshes, or voxels, while the latter two leverage volumetric rendering paradigms to generate photorealistic and continuous scene representations from asynchronous event streams.
@@
4.1 Geometry-based Methods with Traditional Output
Geometry-based methods typically achieve semi-dense or sparse 3D reconstruction in real-time, relying on spatial-temporal event aggregation and geometric priors. Since monocular event cameras cannot directly observe disparities, these methods often require known or estimated motion to compensate for the view angles. Figure 7 and Table 3 provide an overview of these methods. Based on the data acquisition strategy, we categorise these methods into the following two types:
Spatial scanning: Many methods rely on spatial scanning, where the event camera must move through space to accumulate multi-view observations. By capturing asynchronous events from different viewpoints, these methods infer scene structure through motion-induced parallax. In 2016, Kim et al. [41] proposed an approach utilising three interleaved probabilistic filters to estimate camera trajectory, scene log-intensity gradient, and inverse depth. Rebecq et al. [77] introduced EVO, leveraging event projection, edge alignment, and DSI construction for 3D reconstruction. In the same year, Rebecq et al. [75], later in 2018, proposed EMVS, employing event space-sweep and ray density analysis to directly generate a semi-dense 3D depth map, without frame-level data association. A recent study has improved geometry-based methods to achieve dense reconstruction. In 2024, Guan et al. [30] proposed EVI-SAM, a tightly coupled event-image-IMU SLAM framework. It achieves real-time dense 3D reconstruction on a standard CPU, integrating event-based 2D-2D alignment, image-guided depth interpolation, and TSDF fusion [63].
@@ 4.1 Geometry-based Methods with Traditional Output
4.2 Learning-Based Methods with Traditional Output
Deep learning-based methods typically produce non-real-time dense reconstruction. However, traditional RGB image feature extraction techniques cannot be directly applied to raw event data, and image-based event representation is often required as a preprocessing [92]. Table 4 and Figure 7 provide an overview of these methods. Based on the reliance on structured pipelines and external priors, we categorise these methods into the following two types:
@@ 4.2 Learning-Based Methods with Traditional Output
Without pipeline and priors: However, recent methods aim to eliminate the pipeline and priors. In 2023, Chen et al. [12] proposed E2V, which employs a modified ResNet-152 and a U-Net 3D decoder to directly predict dense 3D voxel grids from monocular event frames, achieving event-based 3D reconstruction without external priors. In 2025, Xu et al. [102] extended E2V by introducing a novel event representation, Sobel Event Frame, and an optimal binarisation strategy for event-based 3D reconstruction. By enhancing E2V with Efficient Channel Attention [94], their method significantly improved reconstruction quality.
4.3 Methods with Neural Radiance Fields
@@
Compared to geometry-based or deep learning-based methods, NeRF-based methods avoid explicit feature extraction or geometry construction pipelines. Instead, they rely on end-to-end optimisation guided by photometric or event-based supervision to recover implicit scene structure. NeRF-based methods typically use the Event Accumulate Frame as the main input representation of event streams. Table 5 and Figure 8 provide an overview of these methods. Based on the physical modelling approach, we categorise them into the following two types:
@@
Brightness difference modelling: Some relatively early methods supervise NeRF training only by matching the brightness difference between discrete time points, typically over fixed intervals or adjacent event pairs. These methods do not explicitly model the brightness change rate. In 2023, Hwang et al. [37] proposed Ev-NeRF, which aggregates events through multi-view consistency. By estimating the volumetric density and radiance, the method achieves high-quality depth reconstruction and novel view synthesis. In the same year, Rudnev et al. [81] proposed EventNeRF, which employs event-based volumetric rendering in a self-supervised manner to reconstruct high-quality 3D structures and synthesise new views. Later that year, Klenk et al. [43] proposed E-NeRF, which utilises an event-triggered brightness model along with a no-event loss to enable dense reconstruction. In 2024, Bhattacharya et al. [4] proposed EvDNeRF, the first dynamic event-based NeRF. EvDNeRF uses an event-triggered brightness model and a varied batching strategy to achieve high-fidelity dynamic reconstruction.
Brightness change rate modelling: Some later methods incorporate loss functions that explicitly model the brightness change rate, defined as the brightness difference divided by the corresponding time interval. This enables finer temporal resolution, improves robustness to non-uniform motion, and better aligns with the physical characteristics of event triggering. In 2023, Low et al. [54] proposed Robust e-NeRF. This method introduces a more realistic event generation model and two normalised loss functions: one based on contrast-normalised difference and another on target-normalised temporal gradient. As a result, the model no longer requires known contrast thresholds or explicit event accumulation strategies, enabling robust self-supervised learning. In 2024, Wang et al. [92] proposed Physical Priors Augmented EventNeRF, which incorporates motion and geometric priors and adopts a density-guided patch sampling strategy to enhance structural representation. In 2025, Feng et al. [19] proposed AE-NeRF, which integrates pose correction and hierarchical architecture to reconstruct NeRFs from sparse and asynchronous event streams accurately.
4.4 Methods with 3D Gaussian Splatting
@@
Gaussian Splatting-based methods represent scenes using a set of 3D Gaussian primitives with learnable properties. Each primitive has learnable attributes, including position, covariance, and colour. Similar to NeRF-based methods, most Gaussian Splatting-based methods also use the Event Accumulation Frame as the main input representation of event streams. Table 6 and Figure 9 provide an overview of these methods. Based on the source of supervision and structural guidance, we categorise these methods into the following three types:
Reconstructing images as guidance: Some methods use image frames reconstructed from event streams as a source of visual guidance or as an initial structural prior to facilitate the learning process. In 2024, Zhang et al. [114] proposed Elite-EvGS, which distils prior knowledge from off-the-shelf event-to-video (E2V) models [79, 18]. It uses E2V-generated frames to initialise a coarse 3DGS model and then progressively incorporates raw events to refine scene details through event supervision. In the same year, Yin et al. [106] proposed E-3DGS, which uses a novel temporal-to-intensity mapping, an event-type-specific supervision strategy, and a hybrid optimisation approach. Also in 2024, Yura et al. [110] proposed EventSplat, which combines E2V guided SfM [84] initialisation and spline interpolation. It recovers continuous camera trajectories and achieves high-quality 3D reconstruction.

### v2

@@ III-A Geometry-based Methods
Schraml et al. [32] 2015 Trajectory EAF 3D depth map (✗) Self-collected [32]
Zhu et al. [37] 2018 Velocity Event disparity volume 3D depth map (✓) MVSEC [80]
Zhou et al. [81] 2018 Pose Time surface 3D depth map (✗) MVSEC [80]
@@
Leng et al. [36] 2018 Pose Time surface Point cloud, depth map (✗) Self-collected [36]
Zhu et al. [82] 2019 Pose Event disparity volume 3D depth map (✗) Self-collected [72]
Steffen et al. [72] 2019 - 4D vector Voxel (✗) MVSEC [80]
@@
Zhou et al. [78] 2021 Pose Time surface Point cloud, depth map (✗) MVSEC [80]
Nam et al. [73] 2022 Pose Event Stack 3D depth map (✓) DSEC [83]
@@
Ghosh et al. [71] 2022 Pose Event disparity volume 3D depth map (✗) MVSEC [80], DSEC [83], etc
Ghosh et al. [84] 2024 Pose Event disparity volume 3D depth map (✗) MVSEC [80], DSEC [83], etc
@@
Freitag et al. [79] 2025 Pose Event Stack Point cloud (✓) Self-collected [85]
Hitzges et al. [86] 2025 Pose Event disparity volume 3D depth map (✗) MVSEC [80], DSEC [83]
@@
Without disparity matching: Disparity matching can also be bypassed altogether by directly optimising geometric or temporal consistency across event streams. In 2018, Zhou et al. [81] proposed a novel forward-projection-based depth estimation method, which avoids the need for explicit disparity computation. Instead of finding correspondences between stereo events, their method optimises a temporal consistency energy by projecting candidate depth hypotheses into both stereo time surfaces and measuring their alignment. A coarse-to-fine search strategy is used to obtain semi-dense depth maps in real-time, enabling robust depth reconstruction under challenging conditions. Building upon this idea, Zhou et al. [78] extended the concept to a stereo event-based visual odometry framework. By jointly optimising over spatial and temporal consistency across stereo time surfaces, they estimate not only per-frame depth but also the stereo camera motion. The system runs in real-time and outputs semi-dense depth maps aligned with visual odometry estimates, making it well-suited for 3D reconstruction in dynamic environments. Later in 2022, Ghosh et al. [71] proposed a multi-event-camera fusion framework that leverages Disparity Space Images (DSIs) [87] to accumulate event ray densities from multiple viewpoints. Instead of requiring event correspondences, their method refocuses events in DSI space and performs probabilistic fusion to obtain high-quality depth maps. They further introduce an outlier rejection strategy to enhance robustness and demonstrate generalisation across multiple datasets without fine-tuning. Building upon this idea, in 2024, Ghosh et al. [84] proposed ES-PTAM, which combines a mapping module based on ray density fusion and a tracking module using edge-map alignment, both operating on pure event data. By bypassing disparity matching and processing stereo events in a correspondence-free manner via geometric ray density fusion, the system produces accurate semi-dense depth maps and camera poses in real time. It is a leading stereo visual odometry method that exclusively uses events, without any auxiliary inputs.
III-B Learning-based Methods
@@
Deep learning processing: More recent methods leverage deep neural networks to learn stereo disparity estimation and event representations. In 2019, Zhu et al. [82] proposed an unsupervised deep learning framework, which leverages neural networks to predict depth and motion from stereo event data. They introduce a discretized event volume representation to preserve spatio-temporal structure and apply a motion compensation loss that reduces event blur. In 2022, Nam et al. [73] introduced a deep stereo framework that combines multi-density event stacking with attention-guided encoding via a UNet-ResNet [89, 90] backbone. Their method further incorporates future event prediction during training through a distillation loss [91, 92], enabling real-time, dense depth estimation with high accuracy on the DSEC dataset [93]. In 2025, Hitzges et al. [86] proposed DERD-Net, which estimates depth from event-based ray densities by processing DSIs derived from multi-view event data and known camera poses. By extracting local sub-volumes (Sub-DSIs) and combining 3D convolutions with recurrent units, their method enables efficient, high-resolution depth prediction.
IV Methods with Monocular Event Cameras
@@
Despite these limitations, monocular setups are more lightweight, power-efficient, and easier to deploy and move, making them suitable for embedded systems and mobile equipment. Consequently, monocular event-based 3D reconstruction has attracted increasing research interest and covers a broad range of tasks and 3D representation strategies.
Existing monocular approaches can be broadly divided into four categories based on their underlying modelling techniques and output formats: (IV-A) geometry-based methods with traditional outputs, (IV-B) learning-based methods with traditional outputs, (IV-C) methods with Neural Radiance Fields, and (IV-D) methods with 3D Gaussian Splatting. The first two categories aim to reconstruct explicit 3D outputs such as depth maps, point clouds, meshes, or voxels, while the latter two leverage volumetric rendering paradigms to generate photorealistic and continuous scene representations from asynchronous event streams.
@@
IV-A Geometry-based Methods with Traditional Output
TABLE III: Monocular event camera: Geometry-based methods with traditional outputs
Author Year Priors Event Rep. Real-time Output (Dense?) Dataset
Kim et al. [33] 2016 Trajectory EGA 3D depth map (✗) Self-collected [33]
Rebecq et al. [99] 2016 Trajectory Event-by-event 3D depth map (✗) Self-collected [51]
Rebecq et al. [100] 2016 Pose Event frame 3D depth map (✗) Self-collected [100]
Guan et al. [28] 2024 Trajectory Time surface 3D depth map (✓) DAVIS240C [101]
Elms et al. [102] 2024 Trajectory Time surface Sparse point cloud (✗) TOPSPIN [102]
TABLE IV: Monocular event camera: Learning-based methods with traditional outputs
Author Year Priors Pipeline Event Rep. Model Output Dataset
Baudron et al. [94] 2020 Silhouette EAF E2S(CNN) Mesh (✓) ShapeNet [103]
Xiao et al. [104] 2022 Pose Voxel grid E2VID(RNN-CNN) Mesh (✓) ESIM [105]
Wang et al. [106] 2023 Contour, Trajec. Voxel grid Evac3d(CNN) Mesh (✓) MOEC-3D [106]
Chen et al. [95] 2023 - Event frame E2V(CNN) Voxel (✓) SynthEVox3D [95]
Xu et al. [27] 2025 - Event frame (sobel) E2V(CNN) Voxel (✓) SynthEVox3D [95]
Geometry-based methods typically achieve semi-dense or sparse 3D reconstruction in real-time, relying on spatial-temporal event aggregation and geometric priors. Since monocular event cameras cannot directly observe disparities, these methods often require known or estimated motion to compensate for the view angles. Figure 9 and Table III provide an overview of these methods. Based on the data acquisition strategy, we categorise these methods into the following two types:
Spatial scanning: Many methods rely on spatial scanning, where the event camera must move through space to accumulate multi-view observations. By capturing asynchronous events from different viewpoints, these methods infer scene structure through motion-induced parallax. In 2016, Kim et al. [33] proposed an approach utilising three interleaved probabilistic filters to estimate camera trajectory, scene log-intensity gradient, and inverse depth. In 2016, Rebecq et al. [99] proposed EMVS, employing event space-sweep and ray density analysis to directly generate a semi-dense 3D depth map, without frame-level data association. Later, EMVS was extended into a full SLAM framework as its mapping module, where Rebecq et al. [100] introduced EVO (Event-based Visual Odometry), which integrates event projection, time-surface alignment, and DSI-based depth fusion to jointly estimate camera trajectory and reconstruct 3D geometry in real time. A recent study has improved geometry-based methods to achieve dense reconstruction. In 2024, Guan et al. [28] proposed EVI-SAM, a tightly coupled event-image-IMU SLAM framework. It achieves real-time dense 3D reconstruction on a standard CPU, integrating event-based 2D-2D alignment, image-guided depth interpolation, and TSDF fusion [107].
@@ IV-A Geometry-based Methods with Traditional Output
IV-B Learning-Based Methods with Traditional Output
Deep learning-based methods typically produce non-real-time dense reconstruction. However, traditional RGB image feature extraction techniques cannot be directly applied to raw event data, and image-based event representation is often required as a preprocessing [109]. Table IV and Figure 9 provide an overview of these methods. Based on the reliance on structured pipelines and external priors, we categorise these methods into the following two types:
@@ IV-B Learning-Based Methods with Traditional Output
Without pipeline and priors: However, recent methods aim to eliminate the pipeline and priors. In 2023, Chen et al. [95] proposed E2V, which employs a modified ResNet-152 and a U-Net 3D decoder to directly predict dense 3D voxel grids from monocular event frames, achieving event-based 3D reconstruction without external priors. In 2025, Xu et al. [27] extended E2V by introducing a novel event representation, Sobel Event Frame, and an optimal binarisation strategy for event-based 3D reconstruction. By enhancing E2V with Efficient Channel Attention [118, 119], their method significantly improved reconstruction quality.
IV-C Methods with Neural Radiance Fields
@@
TABLE V: Monocular event camera: NeRF-based Methods
Author Model Yr-Mo Inputs Event Rep. Colourful Dataset
Klenk et al. [35] E-NeRF 2023-01 Event stream EAF ESIM [105]
Hwang et al. [121] Ev-NeRF 2023-03 Event stream EAF IJRR [122], HQF [123]
Rudnev et al. [96] EventNeRF 2023-06 RGB bayer event stream EAF Self-collected [96]
Low et al. [120] Robust e-NeRF 2023-10 Event stream Event-by-event TUM-VIE [124]
Bhattacharya et al. [125] EvDNeRF 2024-01 Event stream EAF Self-collected [125]
Wang et al. [109] NeRF(Enhanced) 2024-05 Event stream EAF PAEv3D [109]
Low et al. [126] Deblur e-NeRF 2024-09 Event stream Event-by-event EDS [127]
Feng et al. [128] AE-NeRF 2025-01 Event stream Event-by-event TUM-VIE [124]
Wang et al. [129] SaENeRF 2025-04 Event stream EAF Rudnev et al. [96]
Figure 11: Visual comparison of several NeRF-based methods using monocular event cameras, adapted from Feng et al. [128], under CC BY-NC-SA 4.0. (a) Result from the monocular deep learning method E2VID combined with NeRF [104]. (b) Ev-NeRF method [121] (c) Robust e-NeRF method [120]. (d) AE-NeRF method [128].
@@
Compared to geometry-based or deep learning-based methods, NeRF-based methods avoid explicit feature extraction or geometry construction pipelines. Instead, they rely on end-to-end optimisation guided by photometric or event-based supervision to recover implicit scene structure. NeRF-based methods typically use the Event Accumulate Frame as the main input representation of event streams and are generally designed for grayscale 3D reconstruction. Particularly, some methods achieve colour 3D reconstruction using only the event stream as input. Low et al. [120, 126] restore colour using gamma correction. Feng et al. [128] adopt a learning-based colour correction method. Wang et al. [129] leverage NeRF’s volume rendering and colour modelling to perform self-supervised colour 3D reconstruction without RGB image supervision. Table V and Figure 10 provide an overview of these methods, and Figure 11 provides a visualisation example of some methods. Based on the physical modelling approach, we categorise them into the following two types:
@@
Brightness difference modelling: Some relatively early methods supervise NeRF training only by matching the brightness difference between discrete time points, typically over fixed intervals or adjacent event pairs. These methods do not explicitly model the brightness change rate. In 2023, Hwang et al. [121] proposed Ev-NeRF, which aggregates events through multi-view consistency. By estimating the volumetric density and radiance, the method achieves high-quality depth reconstruction and novel view synthesis. In the same year, Rudnev et al. [96] proposed EventNeRF, which employs event-based volumetric rendering in a self-supervised manner to reconstruct high-quality 3D structures and synthesise new views. Later that year, Klenk et al. [35] proposed E-NeRF, which utilises an event-triggered brightness model along with a no-event loss to enable dense reconstruction. In 2024, Bhattacharya et al. [125] proposed EvDNeRF, the first dynamic event-based NeRF. EvDNeRF uses an event-triggered brightness model and a varied batching strategy to achieve high-fidelity dynamic reconstruction. In 2025, Wang et al. [129] proposed SaENeRF, which normalizes radiance variations based on accumulated event polarities and introduces zero-event regularization losses, enabling artifact-suppressed and photorealistic 3D reconstruction directly from event streams.
@@
TABLE VI: Monocular event camera: 3DGS-based Methods
Author Model Yr-Mo Inputs Event Rep. Colourful Dataset
Wang et al. [131] EvGGS 2024-05 Event stream Voxel Grid, EAF Ev3DS [131]
Zhang et al. [132] Elite-EvGS 2024-09 Event stream EAF Rudnev et al. [96]
Wu et al. [133] Ev-GS 2024-09 Event stream EGA Self-collected [133]
Huang et al. [134] IncEventGS 2024-10 Event stream EAF TUM-VIE [124]
Yin et al. [130] E-3DGS 2024-10 Event stream Event-by-event EME-3D [130]
Yura et al. [97] EventSplat 2024-12 RGB bayer event stream EAF EDS [127], TUM-VIE [124]
Jeong et al. [135] EOGS 2024-12 Event stream EAF EDS [127]
Zahid et al. [98] E-3DGS 2025-02 RGB bayer event stream EAF E-3DGS dataset [98]
Brightness change rate modelling: Some later methods incorporate loss functions that explicitly model the brightness change rate, defined as the brightness difference divided by the corresponding time interval. This enables finer temporal resolution, improves robustness to non-uniform motion, and better aligns with the physical characteristics of event triggering. In 2023, Low et al. [120] proposed Robust e-NeRF. This method introduces a more realistic event generation model and two normalised loss functions: one based on contrast-normalised difference and another on target-normalised temporal gradient. As a result, the model no longer requires known contrast thresholds or explicit event accumulation strategies, enabling robust self-supervised learning. In 2024, Wang et al. [109] proposed Physical Priors Augmented EventNeRF, which incorporates motion and geometric priors and adopts a density-guided patch sampling strategy to enhance structural representation. Also in 2024, Low et al. [126] proposed Deblur e-NeRF, which models pixel bandwidth to account for event motion blur and introduces a threshold-normalised total variation loss, enabling robust 3D reconstruction directly from motion-blurred event streams. In 2025, Feng et al. [128] proposed AE-NeRF, which integrates pose correction and hierarchical architecture to reconstruct NeRFs from sparse and asynchronous event streams accurately. To enhance visual quality, they use a colour correction network to recover RGB images from log-radiance.
IV-D Methods with 3D Gaussian Splatting
Gaussian Splatting-based methods represent scenes using a set of 3D Gaussian primitives with learnable properties. Each primitive has learnable attributes, including position, covariance, and colour. Similar to NeRF-based methods, most Gaussian Splatting-based methods also use the Event Accumulation Frame as the main input representation of event streams. Table VI and Figure 12 provide an overview of these methods. Based on the source of supervision and structural guidance, we categorise these methods into the following three types:
Reconstructing images as guidance: Some methods use image frames reconstructed from event streams as a source of visual guidance or as an initial structural prior to facilitate the learning process. In 2024, Zhang et al. [132] proposed Elite-EvGS, which distils prior knowledge from off-the-shelf event-to-video (E2V) models [136, 137]. It uses E2V-generated frames to initialise a coarse 3DGS model and then progressively incorporates raw events to refine scene details through event supervision. In the same year, Yin et al. [130] proposed E-3DGS, which uses a novel temporal-to-intensity mapping as visual guidance to facilitate 3D representation learning. The method also incorporates an event-type-specific supervision strategy and a hybrid optimisation approach. Also in 2024, Yura et al. [97] proposed EventSplat, which combines E2V guided SfM [113] initialisation and spline interpolation. It recovers continuous camera trajectories and achieves high-quality 3D reconstruction.